OpenStudy (anonymous):

Identify the mathematical property that describes the following: (3/8)(1)= 3/8 Select one: a. Multiplicative Identity b. Multiplicative Inverse c. Symmetric d. Reflexive

OpenStudy (anonymous):

ill see wat i can do no promisis tho

OpenStudy (anonymous):

okay… so

OpenStudy (anonymous):

In multiplication with real numbers (which this is)...1 is considered the "identity"

OpenStudy (anonymous):

In addition "0" is considered the identity since 3/8 + 0 = 3/8

OpenStudy (anonymous):

:) listen to this guy ^

OpenStudy (anonymous):

So now you know the answer, right?

OpenStudy (anonymous):

haha nice job :)

OpenStudy (anonymous):

Answer is B!!!!!!!!!!!!!!! jk :3

OpenStudy (anonymous):

Haha yeah thanks guys.c: its A.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

The inverse in multiplication is the reciprocal. It's whatever you have to multiply a number by to get the identity...which is 1. For 3/8, the inverse is 8/3 since 3/8 * 8/3 = 1